# VideoWebArena: Evaluating Long Context Multimodal Agents with Video Understanding Web Tasks

### v1

@@ Abstract
Videos are often used to learn or extract the necessary information to complete tasks in ways different than what text and static imagery alone can provide. However, many existing agent benchmarks neglect long-context video understanding, instead focusing on text or static image inputs. To bridge this gap, we introduce VideoWebArena (VideoWA), a benchmark for evaluating the capabilities of long-context multimodal agents for video understanding. VideoWA consists of 2,021 web agent tasks based on manually crafted video tutorials, which total almost four hours of content. For our benchmark, we define a taxonomy of long-context video-based agent tasks with two main areas of focus: skill retention and factual retention. While skill retention tasks evaluate whether an agent can use a given human demonstration to complete a task efficiently, the factual retention task evaluates whether an agent can retrieve instruction-relevant information from a video to complete a task. We find that the best model achieves 13.3% success on factual retention tasks and 45.8% on factual retention QA pairs, far below human performance at 73.9% and 79.3%, respectively. On skill retention tasks, long-context models perform worse with tutorials than without, exhibiting a 5% performance decrease in WebArena tasks and a 10.3% decrease in VisualWebArena tasks. Our work highlights the need to improve the agentic abilities of long-context multimodal models and provides a testbed for future development with long-context video agents.
@@ 1 Introduction
Humans often use videos to complete daily tasks, whether to learn from tutorials or retrieve information from within one or several videos. As we build AI assistants, these multimodal agents must also possess similar capabilities to understand and process videos to accomplish tasks or learn how to accomplish a workflow, plan, and make decisions.
While videos can provide a rich source of information, capturing spatial and temporal dynamics that images or text alone may not convey, integrating video input into multimodal models introduces unique challenges relating to temporal coherence, context retention, or efficient information retrieval over lengthy, extended sequences. These challenges can be further compounded when models are deployed as autonomous agents operating in complex environments. In these scenarios, the ability of models to maintain long-term memory, perform informational retrieval, and adapt to new information continuously is critical for tasks that require sustained engagement over time.
Recent improvements in long-context understanding of large video-capable vision language models (e.g., LLaVaNeXt, LongVILA) have enabled agents to process and understand more information than before, including long video understanding. However, from an evaluative perspective, there remains a significant gap in existing benchmarks that can comprehensively evaluate the agenticn capabilities of these models across diverse multimodal scenarios, particularly those involving video inputs. The requirement for agents to operate across varying modalities and time frames makes developing and properly evaluating long-context multimodal models essential. Existing benchmarks (Wu et al., 2024; Mangalam et al., 2023; Li et al., 2024; Pătrăucean et al., 2023) often fall short in testing for long-term memory retention and multimodal integration within an agent workflow, as they usually focus on only one component. This limits our understanding of how long context multimodal models can generalize and perform in real-world settings as agents.
@@
To address this gap, we present VideoWebArena, a novel, open-source video-based benchmark that evaluates multimodal models’ agentic ability to process, understand, and utilize long-context video inputs to accomplish various tasks. Our contributions can be summarized as follows.
@@
  • Of these 2,021 tasks, VideoWebArena consists of 400 factual retention tasks, which test agents’ abilities to retrieve information from a video to perform tasks, and 1,621 skill retention tasks, which test agents’ abilities to use tutorials in-context to perform tasks more efficiently.
  • We evaluate several video-capable state-of-the-art LLMs, namely GPT-4o and Gemini 1.5 Pro, on our benchmark, providing an overview of these models’ current long-context video understanding capabilities. Our results show that while these models can serve in a limited capacity as video-capable agents, these models are still a far reach from human levels of performance, highlighting a wide gap in the information retrieval and agentic abilities of current state-of-the-art long-context models.
Our code, benchmark, and relevant documentation are available at videowebarena.github.io.
@@ 2 Background
Generally, similar to LLMs, VLM architectures typically revolve around two types—models with either a joint encoder-decoder architecture such as LLaVA and its variations (Liu et al., 2023) or a decoder-only architecture. Encoder-decoder VLMs tend to project different modalities through a shallow neural network or fully connected layer to link modalities. Decoder-style models typically rely on a decoder-only LLM that processes raw inputs (e.g., text tokens, image patches, etc.) such as VILA (Lin et al., 2023) along and its variants such as VILA2 (Fang et al., 2024) and X-VILA (Ye et al., 2024). As multimodal understanding with long-context capability becomes more important in processing more input information like video data, models like LongVILA (Xue et al., 2024) can now process a much larger number of video frames in its input for longer videos.
Agents & Tools.
Recent works have turned to adapting large multimodal models like VLMs into autonomous agents, given their capabilities in visual question answering and high-level reasoning. Many approaches have been further developed to improve or supplement the agentic capabilities of these large models, ranging from data collection and synthesis for specific kinds of training (Lai et al., 2024; Patel et al., 2024) or for use at inference time (Pan et al., 2024; Fu et al., 2024; Koh et al., 2024b; Sarch et al., 2024; Wang et al., 2024). Similar works have shown how certain fine-tuning (Furuta et al., 2024) or prompting techniques (e.g., with Set-of-Marks (Yang et al., 2023), (Chi et al., 2024)) can improve performance in navigating settings like web pages to accomplish tasks. In contrast, others attempt to improve how agents themselves can dynamically compose/search for policies (Sodhi et al., 2024).
@@
As more works adapt LLMs and VLMs into agents, other works have focused on evaluating such efforts. These benchmarks can range across a variety of settings: from general web browsing (Yao et al., 2022; Deng et al., 2023; Zhou et al., 2024; Koh et al., 2024a), where agents are evaluated on their abilities to navigate the web and accomplish specific tasks, to mobile environments, where agents are expected to perform tasks within a mobile OS simulation like Android (Zhang et al., 2024a; Rawles et al., 2024). Other general environments attempt to emulate an OS or computing environment like MMInA (Zhang et al., 2024b), OSWorld (Xie et al., 2024), and Windows Agent Arena (Bonatti et al., 2024) where agents must navigate across multiple computer applications online and offline. Custom environments like WorkArena (Drouin et al., 2024) instead target more specific platforms like ServiceNow in constructing tasks for agent evaluation.
@@
For large multimodal models, long-context capabilities are essential for detailed planning, action, and understanding, especially for the video modality. There have been multiple benchmarks dedicated towards video understanding, with shorter video inputs (Li et al., 2024), around a few minutes, and longer video inputs (Wu et al., 2024; Pătrăucean et al., 2023), up to over an hour. Video understanding benchmarks dedicated to temporal reasoning and thematic reasoning (Xiao et al., 2021; Tapaswi et al., 2016; Lei et al., 2019) also exist. Many of these benchmarks cover subsets and define categories of video understanding tasks related to spatial reasoning, causal reasoning, and temporal reasoning.
@@ 3.1 Summary & Overview
VideoWA centers around six key thematic environments created by VisualWebArena (Koh et al., 2024a) and WebArena (Zhou et al., 2024): Reddit, Classifieds, Shopping, Shopping Admin, Map, and Gitlab. Tables 2 and 2 for a finer characterization of the tasks and videos within the benchmark.
@@ 3.3 Observation Space
The observation space for the VideoWA environment is strictly predicated on the Set-of-Marks observation space in VisualWebArena. The environment uses executable JavaScript code at each step to extract the interactable HTML elements from the webpage and present them in a top-down order. Similarly, the JavaScript code extracts the bounding boxes of each interactable element and a screenshot of the webpage with bounding boxes over the interactable elements is generated for input to the agent along with the text state representation. At each time step, the agent is presented with the overlayed Set-of-Marks screenshot and text observation space, along with the chosen video information to be put into context.
@@ 3.3.1 Action Space
The action space for the agents in the VideoWA environment can be seen in Table 3. The agents are prompted to generate a single action from the action space at each time step. Each action is associated with Playwright Python code that automatically performs the action within the browser. The ‘elem’ parameter in the action represents the unique Set-of-Marks element that can be interacted with from the observation space provided through the environment’s JavaScript code.
@@ 3.4 Task Design
Follow all the repos visited in the video. What are the names of all the visited repos?
@@ 3.5 Video Creation and Skill Retention Tasks
Our benchmark contains 74 unique videos, totaling almost 4 hours of video content (see Table 2 for details)—all of our video tutorials are based on tasks in WebArena and VisualWebArena. We provide our videos online through a YouTube channel and a Google Drive link containing the zip file of all the videos. We formulated these videos by accumulating all the feasible intent templates in WebArena and VisualWebArena. We take 297 unique templates from VisualWebArena and 220 unique templates from WebArena, totaling 1621 total intents. Further details can be found in Appendix A.1
@@ 3.7 Task Evaluation
Each task has an eval and intermediate eval function. We import the automatic functionality from VisualWebArena (Koh et al., 2024a) and WebArena (Zhou et al., 2024) to evaluate our agent tasks. For the intermediate intent evaluation, we use the string-based existing functionality evaluators to assess the agent’s response to the video-based question.
All of our tasks have a final evaluation function (i.e., evaluator) that determines an agent’s reward on each task. The reward is typically binary, returning zero or unity depending on whether the agent performs the task unsuccessfully or successfully. Reward values are determined by evaluating the state of the environment at the end of the agent’s trajectory to determine if said state matches the correct state corresponding to the correct task execution.
@@ 4 Baseline Agents
We evaluate our benchmark using three different types of baseline agents with multimodal models as a backbone. At each step, the agent is given the task objective, 2 in-context examples, current state $s$, and the input video to the objective as context to generate one action.
@@ 4.1 Video In-Context Agent
We define a video input agent that takes the video in at every time step to generate actions. We provide the whole video in-context to the model with the Gemini model. The Gemini model automatically processes the audio, eliminating the need to process audio secretly. The specific prompts we use are in Appendix D. We use Set-of-Marks on the website HTML page, the Set-of-Marks element tree string, and the prompt along with the video as input to the model.
@@ 4.2 Video Frames In-Context Agent
We define a video input agent that takes a set amount of video frames at every time step along with the video audio to generate actions. To obtain the information from the video, we follow the practice from (Wu et al., 2024). We sample 1 frame per second (max 60 frames) for the video and include them into the context for the LLM. In addition, we use OpenAI’s Whisper (Radford et al., 2022) to transcribe the audio and append it to the context. In this way, we can still pass the information from the video to our LLM. This may not be a perfect method as the video information remains missing during framing sampling. However, since most LLMs in the market only support image and text input, it is essential to experiment with this setting. We use GPT-4o, and the prompt can be seen in Appendix D. We again use Set-of-Marks on the website HTML page, the Set-of-Marks element tree string, and the prompt along with 60 video frames and audio transcriptions as input to the model.
@@ 4.3 Video Summary In-Context Agent
We define a summary in-context agent that takes a video summary related to the objective at hand in-context at every time step to generate actions. To obtain this summary, we call GPT-4o and feed the video using 60 frames and the Whisper transcription into the model and prompt it to summarize the video concerning the task at hand. Again, our prompt can be seen in Appendix D. Similarly, the summary agent also uses Set-of-Marks for the observation space and generates actions in the aforementioned action space.
@@ 5 Results
Table 5: Results on VideoWebArena Factual Retention Tasks. Performance of GPT4-o, Gemini 1.5 Pro, and human performance on 400 factual retention tasks broken down by task domain. Final scores indicate the overall task performance (i.e., if the task is completed successfully in its entirety), while intermediate scores measure the performance on the intermediate intents.
@@ 5.1 Model Performance
Visual Perception Task Success Rate 14.1% 11.1% 6.8% 9.3% 7.7%
Audio Perception Task Success Rate 14.8% 18.1% 7.7% 12.5% 11.1%
Full Video Understanding Task Success Rate 15.5% 10.0% 7.2% 10.5% 6.5%
Temporal Reasoning Task Success Rate 13.7% 12.4% 6.2% 10.4% 8.8%
Agentic Easy Task Success Rate 19.5% 12.8% 9.0% 13.0% 8.3%
Agentic Medium Task Success Rate 14.2% 13.4% 5.7% 9.4% 7.7%
Agentic Hard Task Success Rate 10.8% 8.1% 6.2% 9.1% 6.9%
Visual Perception Intermediate Success Rate 32.7 43.9% 43.0% 43.5% 34.0%
Audio Perception Intermediate Success Rate 50.0% 60.2% 62.8% 62.5% 67.9%
Full Video Understanding Intermediate Success Rate 34.2 40.0% 40.9% 41.2% 26.2%
Temporal Reasoning Intermediate Success Rate 35.9 50.5% 50.9% 50.0% 38.9%
Video Easy Intermediate Success Rate 39.5% 52.9% 52.2% 53.2% 47.1%
Video Medium Intermediate Success Rate 39.4% 46.2% 50.4% 48.3% 46.6%
Video Hard Intermediate Success Rate 32.2% 42.4% 40.7% 41.0% 26.1%
Table 7: Factual Retention Results Breakdown: Overall performance breakdown of the baseline agents across all task categories and difficulties in the factual retention set. The summary agent has the best task performance, even without having any visual aspect of the video in context. However, it lags behind in the intermediate VQA intents, as the video frame and video agents all perform very similarly on intermediate tasks.
From Table 5, Table 6 and Table 7, we see varying degrees of agentic performance across the video-capable Gemini and GPT family of models; however, we note several consistent trends across LLM agent results. We comprehensively outline the failure modes in Appendix B. There is no winning baseline agent or model family across skill and factual retention tasks. For factual retention tasks, the summary agent performs the best in task success at 13.3% while the 30 and 100 Frame GPT-4o Agent perform the best in intermediate intent success at 45.8%. For skill retention tasks, we see that long-context models with tutorials actually perform significantly worse than models without tutorials, suggesting that the tutorials introduce negative noise that hurt action selection. Although intermediate scores tend to be higher than final scores, this does not necessarily translate to task success. This is a constant failure mode of the long-context agents, as they can perform the necessary VQA to extract the necessary information for the task at hand but fall short due to hallucinations, action grounding, and high-level planning errors. For example, in Figure 5, the LLM agent successfully identifies the item to buy from the video. Still, it does not successfully plan and complete the intent. We tested on a smaller subset of tasks with the GPT4-o agent and tested on the full set of tasks with the Gemini agent due to compute constraints.

### v2

@@ Abstract
Videos are often used to learn or extract the necessary information to complete tasks in ways different than what text or static imagery can provide. However, many existing agent benchmarks neglect long-context video understanding, instead focusing on text or static image inputs. To bridge this gap, we introduce VideoWebArena (VideoWA), a benchmark for evaluating the capabilities of long-context multimodal agents for video understanding. VideoWA consists of 2,021 web agent tasks based on manually crafted video tutorials, which total almost four hours of content. For our benchmark, we define a taxonomy of long-context video-based agent tasks with two main areas of focus: skill retention and factual retention. While skill retention tasks evaluate whether an agent can use a given human demonstration to complete a task efficiently, the factual retention task evaluates whether an agent can retrieve instruction-relevant information from a video to complete a task. We find that the best model achieves a 13.3% success rate on factual retention tasks and 45.8% on factual retention QA pairs—far below human success rates of 73.9% and 79.3%, respectively. On skill retention tasks, long-context models perform worse with tutorials than without, exhibiting a 5% performance decrease in WebArena tasks and a 10.3% decrease in VisualWebArena tasks. Our work highlights performance gaps in the agentic abilities of long-context multimodal models and provides as a testbed for the future development of long-context video agents.111 Link to code https://github.com/ljang0/videowebarena/
      Link to video https://www.youtube.com/@webarenawarrior or
https://drive.google.com/file/d/17DwmsM7KzBWyz1BN1aq7NHDvgcTIrCgx/view?usp=drive_link.
@@ 1 Introduction
Humans often use videos to complete daily tasks from following tutorials to retrieving information from within one or across several videos. As large foundation models are adapted into AI agents Zheng et al. (2024); Wang et al. (2023); Putta et al. (2024), these multimodal agents may also need similar capabilities to understand and process videos in order to complete various tasks—especially if they need to retrieve and process multimodal information (e.g., from the web) in order to complete said tasks. In these scenarios, the ability of models to maintain long-term memory Zhang et al. (2024b), retrieve and use said information Lewis et al. (2020), and adapt to new information continuously Wu et al. (2024c) is critical for complex, multi-step tasks that require sustained engagement over time.
While videos contain spatial and temporal dynamics that images or text alone may not convey, integrating video input into multimodal models introduces unique challenges (Tang et al., 2024) such as temporal coherence, context retention, or efficient retrieval over lengthy sequences. However, recent advancements in long-context understanding of large video-capable vision language models, such as LLaVaNeXt (Liu et al., 2023), LongVILA (Xue et al., 2024), and Gemini (Google, 2023) have enabled these agentic models to not only process and understand more information than before, including long video understanding, but also generate and act more freely than before. While many existing agentic benchmarks test, they focus primarily on text and image modalities for tasks like interactive QA Mialon et al. (2023), computer tasks Bonatti et al. (2024); Deng et al. (2023), etc.
As a result, from an evaluative perspective, there remains a significant gap in existing benchmarks that can comprehensively evaluate the agentic capabilities of these models across diverse multimodal scenarios, particularly those involving video inputs. The requirement for agents to operate across varying modalities and time frames makes developing and properly evaluating long-context multimodal models essential. Existing benchmarks (Wu et al., 2024b; Mangalam et al., 2023; Li et al., 2024; Pătrăucean et al., 2023) often fall short in testing for long-term memory retention and multimodal integration within an agentic workflow, limiting our understanding of how long context multimodal models can generalize and perform in real-world settings as agents.
@@
To address this gap, we present VideoWebArena, a novel, open-source video-based benchmark that evaluates multimodal models’ agentic ability to process, understand, and utilize long-context video inputs to accomplish various tasks. We summarize our contributions as follows.
@@
  • Of these tasks, VideoWebArena consists of 400 factual retention tasks, which test agents’ abilities to retrieve information from a video to perform tasks, and 1,621 skill retention tasks, which test agents’ abilities to use tutorials in-context to perform tasks more efficiently.
  • We evaluate popular and recently video-capable state-of-the-art LLMs (e.g., GPT-4o and Gemini 1.5 Pro) on our benchmark to provide an overview of these models’ current long-context video understanding capabilities. Our results show that these models can serve as video-capable agents, but in a limited capacity as they are still far from human levels of performance, highlighting a considerable gap in the information retrieval and agentic abilities of current state-of-the-art long-context models.
For details on the benchmark’s code and documentation, please see our reproducibility statement at the end of this paper.
@@ 2 Background
Benchmark Duration (Avg) Multi-domain? Multi-skill? Audio? Agentic?
MSVD-QA (Xu et al. (2017)) 0.33
MSRVTT-QA (Xu et al. (2017)) 0.25
ActivityNet-QA (Yu et al. (2019)) 1.85
NExT-QA (Xiao et al. (2021a)) 0.73
MoVQA (Zhang et al. (2023)) 16.5
MovieChat-1K (Song et al. (2023)) 9.4
MVBench (Li et al. (2023)) 0.27
PerceptionTest (Pătrăucean et al. (2023)) 0.38
CinePile (Rawal et al. (2024)) 2.7
EgoSchema (Mangalam et al. (2024)) 2.7
LongVideoBench (Wu et al. (2024a)) 7.9
Video-MME (Fu et al. (2024a)) 16.9
VideoWebArena (this work) 3.1
Table 1: Comparison of relevant benchmarks/datasets focusing on video modality understanding. Duration denotes the average video duration in minutes, multi-domain denotes if videos cover a diverse range across multiple domains, multi-skill denotes whether the benchmark tests multiple types/aspects of video understanding, audio denotes if the video’s audio information is included or used, and agentic denotes whether the benchmark uses video QA as part of an agent’s workflow as a tutorial/aid, an input required for completing a multi-step task, etc. We also note that, unlike many benchmarks here, VideoWebArena comes with a fully interactive environment, uses open-answer questions relating to video understanding (rather than strictly providing multiple choice options to the model), evaluates task completion on both the final state but also intermediary states, tests across five types/aspects of video understanding grouped into two categories (factual vs. skill retention) and more.
@@
Generally, similar to LLMs, VLM architectures typically revolve around two types—models with either a joint encoder-decoder architecture such as LLaVA and its variations (Liu et al., 2023) or a decoder-only architecture. Encoder-decoder VLMs tend to project different modalities through a shallow neural network or fully connected layer to link modalities. Decoder-style models typically rely on a decoder-only LLM that processes raw inputs (e.g., text tokens, image patches, etc.) such as VILA (Lin et al., 2023) and its variants like VILA2 (Fang et al., 2024) and X-VILA (Ye et al., 2024). As multimodal understanding with long-context capability becomes more important in processing more input information like video data, models like LongVILA (Xue et al., 2024) can now process a much larger number of video frames in its input for longer videos.
Agents.
Recent works have turned to adapting large multimodal models like VLMs into autonomous agents, given their capabilities in visual question answering and high-level reasoning. Many approaches have been further developed to improve or supplement the agentic capabilities of these large models, ranging from data collection and synthesis for specific kinds of training (Lai et al., 2024; Patel et al., 2024) or for use at inference time (Pan et al., 2024; Fu et al., 2024b; Koh et al., 2024b; Sarch et al., 2024; Wang et al., 2024). Similar works have shown how certain fine-tuning (Furuta et al., 2024) or prompting techniques (e.g., Set-of-Marks (Yang et al., 2023), (Chi et al., 2024)) can improve performance in navigating settings like web pages to accomplish tasks. Others attempt to improve how agents themselves dynamically compose/search for policies (Sodhi et al., 2024).
@@
As more works adapt LLMs and VLMs into agents, other works have focused on evaluating such efforts. These benchmarks can range across a variety of settings: from general web browsing (Yao et al., 2022; Deng et al., 2023; Zhou et al., 2024; Koh et al., 2024a), where agents are evaluated on their abilities to navigate the web and accomplish specific tasks, to mobile environments, where agents are expected to perform tasks within a mobile OS simulation like Android (Zhang et al., 2024a; Rawles et al., 2024). Other general environments attempt to emulate an OS or computing environment like MMInA (Zhang et al., 2024c), OSWorld (Xie et al., 2024), and Windows Agent Arena (Bonatti et al., 2024) where agents must navigate across multiple computer applications online and offline. Custom environments like WorkArena (Drouin et al., 2024) instead target more specific platforms like ServiceNow in constructing tasks for agent evaluation.
@@
For large multimodal models, long-context capabilities are essential for detailed planning, action, and understanding, especially for the video modality. There have been multiple benchmarks dedicated towards video understanding, with shorter video inputs (Li et al., 2024), around a few minutes, and longer video inputs (Wu et al., 2024b; Pătrăucean et al., 2023), up to over an hour. Video understanding benchmarks dedicated to temporal reasoning and thematic reasoning (Xiao et al., 2021b; Tapaswi et al., 2016; Lei et al., 2019) also exist. Many of these benchmarks cover subsets and define categories of video understanding tasks related to spatial reasoning, causal reasoning, and temporal reasoning.
@@ 3.1 Summary & Overview
VideoWA centers around six key thematic environments created by VisualWebArena (Koh et al., 2024a) and WebArena (Zhou et al., 2024): Reddit, Classifieds, Shopping, Shopping Admin, Map, and Gitlab. See Tables 3 and 3 for a finer characterization of the tasks and videos within the benchmark.
@@ 3.3 Observation Space
The observation space for our VideoWA environment is predicated on the Set-of-Marks observation space in VisualWebArena. The environment uses executable JavaScript code at each step to extract the interactable HTML elements from the webpage and present them in a top-down order. Similarly, the JavaScript code extracts the bounding boxes of each interactive element and a screenshot of the webpage with bounding boxes over said elements is generated for input into the agent along with the text state representation. At each step, the agent is presented with the overlaid Set-of-Marks screenshot and text observation space, along with the chosen video information to be put into context.
@@ 3.3.1 Action Space
The action space for the agents in the VideoWA environment can be seen in Table 4. Agents are prompted to generate a single action from the action space at each time step. Each action is associated with Playwright Python code that automatically performs the action within the browser. The ‘[elem]’ represents the unique Set-of-Marks element that can be interacted with from the observation space provided through the environment’s JavaScript code.
@@ 3.4 Task Design
Can you assign the issue with a critical priority label that showed up earliest in the video to @earlev4? What was the the name of the issue with a critical priority label that showed up earliest in the video?
@@ 3.5 Video Creation and Skill Retention Tasks
Our benchmark contains 74 unique videos, totaling almost 4 hours of video content (see Table 3 for details)—all of our video tutorials are based on tasks in WebArena and VisualWebArena. We provide our videos online through a YouTube channel and a Google Drive link containing the zip file of all the videos. We formulated these videos by accumulating all the feasible intent templates in WebArena and VisualWebArena. We take 297 unique templates from VisualWebArena and 220 unique templates from WebArena, totaling 1,621 total intents. Further details can be found in Appendix A.1
@@ 3.7 Task Evaluation
Each task has an ‘eval’ and intermediate ‘eval’ function. We import the automatic functionality from VisualWebArena (Koh et al., 2024a) and WebArena (Zhou et al., 2024) to evaluate our agent tasks. For the intermediate intent evaluation, we use the string-based existing functionality evaluators to assess the agent’s response to the video-based question.
All tasks have a final evaluation function (i.e., evaluator) that determines an agent’s reward on each task. The reward is typically binary, returning zero or unity depending on whether the agent performs the task unsuccessfully or successfully. Reward values are determined by evaluating the state of the environment at the end of the agent’s trajectory to determine if said state matches the correct state corresponding to the correct task execution.
@@ 4 Baseline Agents
We evaluate our benchmark using three different types of baseline agents with multimodal models as a backbone. Each type is distinguished by the type of video input provided to the model/agent. At each step, the agent is given the task objective, 2 in-context examples, current state $s$, and the input video to the objective as context to generate one action. We describe each type in detail below.
@@ 4.1 Video In-Context Agent
We define a video input agent that takes the video in at every time step to generate actions. We provide the whole video in-context to the model with the Gemini model. The Gemini model automatically processes the audio, eliminating the need to process audio separately. The specific prompts we use are in Appendix E. We use Set-of-Marks on the website HTML page, the Set-of-Marks element tree string, and the prompt along with the video as input to the model.
@@ 4.2 Video Frames In-Context Agent
We define a video input agent that takes a set amount of video frames at every time step along with the video audio to generate actions. To obtain the information from the video, we follow the practice from (Wu et al., 2024b). We sample 1 frame per second (max 60 frames) for the video and include them into the context for the LLM. In addition, we use OpenAI’s Whisper (Radford et al., 2022) to transcribe the audio and append it to the context. In this way, we can still pass the information from the video to our LLM. This may not be a perfect method as the video information remains missing during framing sampling. However, since most LLMs on the market only support image and text input, it is essential to experiment with this setting. We use GPT-4o, and the prompt can be seen in Appendix E. We again use Set-of-Marks on the website HTML page, the Set-of-Marks element tree string, and the prompt along with 60 video frames and audio transcriptions as input to the model.
@@ 4.3 Video Summary In-Context Agent
We define a summary in-context agent that takes a video summary related to the objective at hand in-context at every time step to generate actions. To obtain this summary, we call GPT-4o and feed the video using 60 frames and the Whisper transcription into the model and prompt it to summarize the video concerning the task at hand. Again, our prompt can be seen in Appendix E. Similarly, the summary agent also uses Set-of-Marks for the observation space and generates actions in the aforementioned action space.
@@ 5 Results
Phi-3.5V (30 Frames) Classifieds 0.0% 8.3% 8.2
Gitlab 2.9% 10.0% 6.9
Map 0.0% 26.7% 16.9
Reddit 0.0% 14.1% 7.5
Shopping (admin) 2.1% 12.8% 12.2
Shopping 0.08% 5.8% 8.6
Total 1.2% 13.0% 10.1
@@
Table 6: Results on VideoWebArena Factual Retention Tasks. Performance of GPT4-o, Gemini 1.5 Pro, Phi-3.5V, and human performance on 400 factual retention tasks broken down by task domain. Final scores indicate the overall task performance (i.e., if the task is completed successfully in its entirety), while intermediate scores measure the performance on the intermediate intents.
@@ 5.1 Model Performance
Phi-3.5V
(30 Frames)
Visual Perception Task Success Rate 14.1% 11.1% 6.8% 9.3% 7.7% 0.8%
Audio Perception Task Success Rate 14.8% 18.1% 7.7% 12.5% 11.1% 1.1%
Full Video Understanding Task Success Rate 15.5% 10.0% 7.2% 10.5% 6.5% 1.1%
Temporal Reasoning Task Success Rate 13.7% 12.4% 6.2% 10.4% 8.8% 0.5%
Agentic Easy Task Success Rate 19.5% 12.8% 9.0% 13.0% 8.3% 0.0%
Agentic Medium Task Success Rate 14.2% 13.4% 5.7% 9.4% 7.7% 1.6%
Agentic Hard Task Success Rate 10.8% 8.1% 6.2% 9.1% 6.9% 0.7%
Visual Perception Intermediate Success Rate 32.7% 43.9% 43.0% 43.5% 34.0% 7.9%
Audio Perception Intermediate Success Rate 50.0% 60.2% 62.8% 62.5% 67.9% 23.5%
Full Video Understanding Intermediate Success Rate 34.2% 40.0% 40.9% 41.2% 26.2% 7.3%
Temporal Reasoning Intermediate Success Rate 35.9% 50.5% 50.9% 50.0% 38.9% 8.0%
Video Easy Intermediate Success Rate 39.5% 52.9% 52.2% 53.2% 47.1% 17.3%
Video Medium Intermediate Success Rate 39.4% 46.2% 50.4% 48.3% 46.6% 9.0%
Video Hard Intermediate Success Rate 32.2% 42.4% 40.7% 41.0% 26.1% 5.3%
Table 8: Factual Retention Results Breakdown: Performance breakdown of baseline agents across all task categories and difficulties in the factual retention set. The summary agent has the best task performance, even without having any visual aspect of the video in context. However, it lags behind in the intermediate VQA intents, as the video frame and video agents all perform very similarly on intermediate tasks.
From Tables 6, 7, and 8, we see varying degrees of agentic performance across the video-capable Gemini and GPT family of models; however, we note several consistent trends across LLM agent results. We outline failure modes in Appendix B. There is no best performing baseline agent or model across skill and factual retention tasks. For factual retention tasks, the summary agent performs the best in task success at 13.3% while the 30 and 100 Frame GPT-4o Agent perform the best in intermediate intent success at 45.8%. For skill retention, we see that long-context models with tutorials actually perform significantly worse than models without tutorials, suggesting that tutorials introduce negative noise that hurt action selection. Although intermediate scores tend to be higher than final scores, this did not translate to task success. This is a constant failure mode of the long-context agents, as they can perform the necessary VQA to extract the necessary information for the task at hand but fall short due to hallucinations, action grounding, and high-level planning errors. For example, in Figure 5, the LLM agent successfully identifies the item to buy from the video. Still, it does not successfully plan and complete the intent. We tested on a smaller subset of tasks with the GPT4-o agent and tested on the full set of tasks with the Gemini agent due to compute constraints.